By measuring the amount of time it takes a component of a product to move from one workstation to the next, an engineer has estimated that the standard deviation is 3.18 seconds.
(A) How many measurements should be made in order to be 98% certain that the maximum error of estimation will not exceed 2.0 seconds?

(B) What sample size is required for a maximum error of 1.0 seconds
Please walk me through this

Answer by stanbon(75887)

the standard deviation is 3.18 seconds.
(A) How many measurements should be made in order to be 98% certain that the maximum error of estimation will not exceed 2.0 seconds?
---
Since E = z*s/sqrt(n),
n = [z*s/E]^2
----
Also, 98% confidence has a left tail of 1% and a right tail of 1%
-invNorm(0.01) = 2.3263
------
Your Problem:
n = [2.3263*3.18/2]^2 = 14 when rounded up
=====================================================
(B) What sample size is required for a maximum error of 1.0 seconds
n = [2.3263*3.18/1]^2 = 55 when rounded up
=============
Cheers,
Stan H.
===============